OpenStudy (anonymous):

9. What is the x-intercept of this equation? 2x + 4y = 4 A. 1 B. 2 C. 3 D. 4 10. What is the y-intercept of this equation? 2x + 4y = 4 A. 1 B. 2 C. 3 D. 4

OpenStudy (campbell_st):

let y = 0 and solve for x and for the y-intercept... let x = 0 and solve for y

OpenStudy (anonymous):

9. B 10.D ?

OpenStudy (campbell_st):

I think your y- intercept is wrong 2x + 4y = 4 x = 0 so 4y = 4 solve for y

OpenStudy (anonymous):

9 is right then? and im still confused sorrry

OpenStudy (anonymous):

2x0=0

OpenStudy (anonymous):

A?

OpenStudy (campbell_st):

ok... a y-itercept occures when x = 0 so if you substitute it into the equation 2x + 4y = 4 the you have 4y = 4 to solve for y... divide both sides by 4...

OpenStudy (anonymous):

4/2=2 4/4=1

OpenStudy (anonymous):

B and A

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

thank you so much!